CAT Chat Minutes

August 4, 2000

 

Information and Follow-up Items

1. The TV Screen image in LOM 438 is not clear - can anything be done to fix this?

The image is fuzzy because the signal has deteriorated due to additional taps being added to the video distribution system. Additional amplifiers are being added and that will improve the signal.

2. Why must the ring current be at 100 mA before beam can be deemed useable again?

All of the saved files from the last acceptable orbit were taken at the 100 mA level, as defined by the mechanical position of the BPMs under that thermal load. However, we are reviewing the level of position error that would exist at lower currents.

 

User Information

On Wednesday afternoon, August 2, in one of the straight sections housing the rf cavities, a roughting pump backing up a turbo pump failed. Despite the interlock system and before the isolation valves could fully close, a burst of air got into the rf cavities and the surrounding girders. The bake out of the affected girder area was completed by Thursday afternoon, August 3rd.

The conditioning of the rf cavities is a more complicated process. We are attempting to condition the rf cavities are rapidly as possible without compromising the rf system. We are conditioning with reduced beam current, and due to the low rf voltage, the lifetime is low. Gap voltages and the beam current will be increased as the vacuum in the rf cavities continues to improve.

There is much discussion going on about when to return the beam to user operations status. Currently, the shutter permission has been given and the orbit has been placed where it should approximately be. This time is not being counted as "user hours" of beam time. Users are asked to not request any sterring until 100 mA is reached. The user community will be kept informed about plans and timetables.